OpenStudy (anonymous):

How do find the exact value of arctan(tan -pi/8)?

rishavraj (rishavraj):

the same applies here....

OpenStudy (anonymous):

Really? How many trigonometric functions does it apply to? If I may ask.

rishavraj (rishavraj):

hold on

OpenStudy (anonymous):

Okay, thank you so much.

rishavraj (rishavraj):

check out this one

OpenStudy (anonymous):

Thank you! And if I may ask a quick question. What do you do when the function is something like cos(arcsin)?

rishavraj (rishavraj):

cos(arc sin x) = sin(arc cos x) = sqr root (1-x^2) and

rishavraj (rishavraj):

\[1\le x \le 1\]

OpenStudy (anonymous):

Thank you! I really appreciate all the help!

rishavraj (rishavraj):

anytime!!!